The following equations represent the inverse supply and demand functions in the market for Good A: PC = 80 - ½ QD PP = 14 + QS where PC and PP are the prices paid by consumers and received by producers respectively. QD and QS are the quantities demanded and supplied, respectively.
Suppose the government is considering imposing a tax of $6 per unit of Good A.
a) (2pts) Compute the competitive market equilibrium price and output without the tax.
b) (4pts) Compute producer surplus and consumer surplus without the tax.
Answer #1

ÀQue) Page 1 GoodA whose demand cu ve 169 PC = 80-ID d curve is given as Taking PC-O PC-80 D=160 AD-O LOD = 80-PC QD = 160-2PCPoice Supply curve dd Cwe Od Qty 10 20 30 409450 G 70 80 90 100 110 120 130 140 150 160170 Producer Surplus -> Asea above suppage-2 with Tax = $6 perunit Now stewillt Le 9+11+ Gewriting demand and Supply Curve QD = 160-2P and OS = D-14 At equilibriumRest; C60-50 x 40 = 6X 40 =$240 is will be the government onevenue Now Producer Nweplus = I x(54-14) x 40 =+*40x 40 Producer